Once again, I could use your assistance guys (I will be needing a lot of it until the second week of January when my school term is over).
(-3 ____) is a solution of y = 4 - 3x
Special thanks to mananth for giving the right solution to the complicated "pi" question last week (to solve for h not r), you're really good!
Thanks in advance.
Monica

Answer by mananth(16946)

(-3 ____) is a solution of y = 4 - 3x
...
y = 4 - 3x
plug value of -3 in the equation
y = 4-3*(-3)
y = 4 +9
y =13
..
(-3,13)
..
m.ananth@hotmail.ca